OpenStudy (anonymous):

How do I put 5y-x=10 in slope intercept form

OpenStudy (anonymous):

Answer: y = -(1/5)x - 2 Steps: The y-intercept form is y = mx + c We need to convert -x -5y=10 to that form. 1. Add x to bot sides -5y = x + 10 2. Divide both sides by -5 y = -(1/5)x - 2